media%2F283%2F283a14d9-12e2-4f0a-bbe5-24
Answer #1

Given

Wavelength is Lambda = 0.300 nm

angle for the first maximum is theta = 24.5 degrees

spacing between the planes of the crystal is d =?

Part A

From Bragg's law 2d sin theta = n *lambda

d = n*lambda/(2* sin theta)

for n =1

d = (1*0.3*10^-9)/(2*sin24.5) m

d = 3.617131522*10^-10 m

d = 0.3617131522 nm

Part B

for second maximum

d = n*lambda/(2* sin theta)

for n =2

d = (2*0.3*10^-9)/(2*sin24.5) m

d = 7.2342630*10^-10 m

d = 0.72342630 nm